OpenStudy (anonymous):

y+(-6)=-9 can anyone help me out?

OpenStudy (anonymous):

y = -9 - (-6) y = -3

OpenStudy (compassionate):

Asad0000, it is best to lead the member to the answer.

OpenStudy (anonymous):

these ar so confuseing

OpenStudy (anonymous):

In a case as simple as this, I think the best thing was to give the answer.

OpenStudy (compassionate):

Kellyp, what don't you understand?

OpenStudy (anonymous):

how do you write an equasion?

OpenStudy (anonymous):

i have a hard time learning math dont unerstand it

OpenStudy (compassionate):

An equation has an equal sign in it. x = 1<-- equation x + y = 2 <-- equation 2y - 5 = 9 <-- equation

OpenStudy (anonymous):

thirteen multiplied by h is onehundred and four

OpenStudy (compassionate):

lolwut

OpenStudy (anonymous):

i have to write it as an equasion

OpenStudy (compassionate):

13 x h = ??? What does 13 times h equal? 13 x y = (Put the number here)

OpenStudy (anonymous):

thats were im haveing issues what h means

OpenStudy (compassionate):

h is a variable. Basically, h can be any number. For example, 13 x h = 103 Just replace h with any number you want 13 x 1 = 103 13 = 103, this is false because 13 does not equal 103.

OpenStudy (anonymous):

really its very confuseing

OpenStudy (compassionate):

You're having troubles with letters, right?

OpenStudy (anonymous):

yes i do not understand this kind of math.. i am trying to learn it but it is hard

OpenStudy (compassionate):

This type of math is called algebra. It is very simple. Please, view this link to understand what these, "letters," mean. They can be confusing at first, but they're actually very easy to understand. http://www.mathsisfun.com/definitions/variable.html

OpenStudy (compassionate):

Also, please click this and watch the video. It is sure to help you understand algebra! http://www.brightstorm.com/math/algebra/pre-algebra/defining-variables/

OpenStudy (anonymous):

im gona look at it i quite school when i was young and quite before i learned it so now im havein a hard time

OpenStudy (compassionate):

You'll do fine. This type of math is really easy to get a hang of. Please look at both the links I posted.

OpenStudy (anonymous):

i am def gona

OpenStudy (compassionate):

Let me know once you've viewed those links and if you understand or not.

OpenStudy (anonymous):

ok the one that i was wondering about thirteen multiplied by h is one hundred and four can u explain that one

OpenStudy (compassionate):

Okay, so lets focus on the first part of this word problem. It's saying thirteen is being multiplied by h. Basically, h is a variable. H can be any number, we don't know what h is currently, but we do know it's being multiplied by h. So, "Thirteen multiplied by h," would be: 13 x h <-- Now this is equal to 104. 13 x h = 104 <-- Because the equation tells us 13 times h is equal to 104.

OpenStudy (anonymous):

how do you come up with h is equal to 104 ?

OpenStudy (compassionate):

Because, " thirteen multiplied by h is EQUAL TO ONE HUNDRED AND FOUR."

OpenStudy (anonymous):

so do i have to find a number that is multiplied by 13 to make 104

OpenStudy (compassionate):

Yes! :D You're getting it!

OpenStudy (compassionate):

13 x h = 104 13 x 1 = 13 <-- this is not 104 13 x 2 = 26 <-- this is not 104.

OpenStudy (anonymous):

ok so would it be 8

OpenStudy (compassionate):

Yes! Because 13 x 8 = 104!!! :D

OpenStudy (anonymous):

because 13 x 8 1s 104

OpenStudy (compassionate):

Mhm. You see, it's easy! Try this one: 5 x y = 25

OpenStudy (anonymous):

ok 5x5=25

OpenStudy (compassionate):

Yes! So y is just any number that can be multiplied by 5 to equal 25.

OpenStudy (anonymous):

ok here is one y+(-6)=-9

OpenStudy (compassionate):

Do you know the first step?

OpenStudy (anonymous):

what is the - signs for

OpenStudy (anonymous):

is the y any number?

OpenStudy (compassionate):

Okay. You need to know what pemdas is. Watch this video. https://www.khanacademy.org/math/arithmetic/multiplication-division/order_of_operations/v/introduction-to-order-of-operations

OpenStudy (anonymous):

is this algebra? 5(b+8)

OpenStudy (compassionate):

Anything with letters is. Sorry, but I have to go. Catch you on the flip side.

OpenStudy (anonymous):

ok thank you so much for your help